# ABS-Mamba: SAM2-Driven Bidirectional Spiral Mamba Network for Medical Image Translation

### v1

@@ 2.1 Over of ABS-Mamba
ABS-Mamba is a novel architecture that synergistically integrates SAM2’s vision transformer with CNNs-based local feature extractors, enhanced by the recent Mamba module for longitudinal context modeling. As illustrated in Fig. 1, the hybrid encoder employs a frozen SAM2 image encoder with 16$\times$16 patch embedding to process input image with high-resolution of 512 $\times$ 512, while the CNNs branch extracts hierarchical features from images with low-resolution of 256$\times$256. The RFFN dynamically fuses multi-scale features, enabling the model to learn a unified representation that harmonizes complementary information from dual-resolution medical imaging protocols. During training, we adapt Low-Rank Adaptation (LoRA+) on only 19.47% of SAM2’s decoder parameters, maintaining the pretrained model’s diagnostic generalization and enabling domain-specific refinement [22].
@@ 2.2.1 Robust Feature Fusion Network(RFFN)
{$F_{\text{even}}&=F_{l}[:,2k::2,:,:]\\ F_{\text{odd}}&=F_{l}[:,2k+1::2,:,:]\end{aligned}\right.\quad\forall k\in\{0,1,\dots,\left\lfloor C_{in}/2\right\rfloor-1\}$ (1)
where $F_{even},F_{odd}\in\mathbb{R}^{B\times(C_{in}/2)\times H\times W}$. Guided by the feature pyramid theory, convolutional kernels of differing scales (3$\times$3 and 5$\times$5) are implemented through depth-wise separable convolutions on the decoupled subfeatures respectively, enhancing cross-scale feature completeness [23]. After multi-kernel processing, the decomposed features undergo:
@@
where $\alpha$ = 0.05 governs the negative signal retention ratio, enhancing robust-ness against adversarial samples. $\mathcal{W}_{\textsl{pw}}$ denotes the 1$\times$1 pointwise convolution.
@@ 2.2.2 Cross-Resolution Hierarchical Feature Fusion
: Let $f_{i}^{a}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ represent the output feature from the $i$-th layer of CNNs Encoder.The SAM2-Hiera encoder outputs four hierarchical $f_{j}^{b}\in\mathbb{R}^{C_{j}\times H_{j}\times W_{j}}$. Where $C_{i},C_{j}\in\{144,288,576,1152\}$ [24]. Let $f_{ij}^{ab}$ denotes the fused feature between the $i$-th and $j$-th layers via RFFN$(f_{i}^{a},f_{j}^{b})$.
The first layer of the encoder processes single-channel input $X_{a}\in\mathbb{R}^{1\times 256\times 256}$, preserving resolution while expanding the channel dimension to 3. Reflection padding(nn.ReflectionPad2d(3))extends it to $262\times 262$, mirroring boundaries to reduce artifacts.A $7\times 7$ convolution (a stride of 2) outputs a $3\times 256\times 256$ feature map, followed by BatchNorm and ReLU, yielding $f_{1}^{a}\in\mathbb{R}^{3\times 256\times 256}$. For the second layer, a 3$\times$3 convolution with stride=2 downsamples $f_{1}^{a}$ to $f_{2}^{a}\in\mathbb{R}^{144\times 128\times 128}$. After normalization and activation procedures, the RFFN integrates $f_{2}^{a}$ and $f_{1}^{b}$ via residual connections, generating the fused feature $f_{21}^{ab}\in\mathbb{R}^{144\times 128\times 128}$. The third layer follows this paradigm, further downsampling to $f_{3}^{ab}\in\mathbb{R}^{288\times 64\times 64}$ and performing secondary fusion with $f_{2}^{b}$, ultimately producing $f_{32}^{ab}\in\mathbb{R}^{288\times 64\times 64}$.
@@ 2.3 Bidirectional Mamba Residual Network
$\hat{X}_{1}&=\text{Reshape}(X)\in\mathbb{R}^{B\times C\times L}\\ \tilde{X}_{1}&=S\cdot\hat{X}_{1}^{\top}\in\mathbb{R}^{B\times L\times C}\end{aligned}\right.L=H\times W$ (3)
@@
{$Y_{1}&=\text{Mamba}_{1}(\text{LN}(\tilde{X}_{1}))\\ Y_{2}&=\text{Mamba}_{2}(\text{LN}(\tilde{X_{2}}))\end{aligned}\right.\quad\text{Mamba}(Z)=\sigma(W_{\text{conv}}*Z)\odot(W_{\text{proj}}\cdot\text{SSM}(Z))$ (4)
@@ 3.2 Results
Table 1: Quantitative comparison of CT to MRI conversion task on Brain and Pelvis against other baseline methods. PSNR (dB) and SSIM are listed as mean±$\pm$±std across the test set. The best results are highlighted in bold.
@@
$\text{ResViT}_{\text{(TMI'22)}}$ 0.785$\pm$0.012 21.58$\pm$0.85 0.684$\pm$0.022 24.65$\pm$1.01
$\text{TransUNet}_{\text{(MIA'24)}}$ 0.842$\pm$0.011 24.47$\pm$0.79 0.728$\pm$0.024 26.32$\pm$0.99
[…]
$\text{VMamba}_{\text{(ICML'24)}}$ 0.859$\pm$0.014 25.23$\pm$1.01 0.736$\pm$0.022 26.49$\pm$0.97
$\text{DiffMa}_{\text{(arXiv'24)}}$ 0.866$\pm$0.009 25.84$\pm$0.88 0.743$\pm$0.012 27.07$\pm$1.01
@@
Table 2: Quantitative comparison of T1 to T2 conversion task on Brain against other baseline methods. PSNR (dB) and SSIM are listed as mean±$\pm$±std across the test set. The best results are highlighted in bold.
@@
$\text{ResViT}_{\text{(TMI'22)}}$ 0.865$\pm$0.013 24.58$\pm$0.75
$\text{TransUNet}_{\text{(MIA'24)}}$ 0.882$\pm$0.012 26.97$\pm$0.69
$\text{VMamba}_{\text{(ICML'24)}}$ 0.903$\pm$0.008 28.42$\pm$0.72

### v2

@@ 2.1 Over of ABS-Mamba
ABS-Mamba is a novel architecture that synergistically integrates SAM2’s vision transformer with CNNs-based local feature extractors, enhanced by the recent Mamba module for longitudinal context modeling. As illustrated in Fig. 1, the hybrid encoder employs a frozen SAM2 image encoder with $\times16$ patch embedding to process input image with high-resolution of 512 $\times$ 512, while the CNNs branch extracts hierarchical features from images with low-resolution of $\times256$. The RFFN dynamically fuses multi-scale features, enabling the model to learn a unified representation that harmonizes complementary information from dual-resolution medical imaging protocols. During training, we adapt Low-Rank Adaptation (LoRA+) on only 19.47% of SAM2’s decoder parameters, maintaining the pretrained model’s diagnostic generalization and enabling domain-specific refinement [22].
@@ 2.2.1 Robust Feature Fusion Network(RFFN)
$\left\{\begin{aligned} F_{\text{even}}&=F_{l}[:,2k::2,:,:]\\ F_{\text{odd}}&=F_{l}[:,2k+1::2,:,:]\end{aligned}\right.\quad\forall k\in\{0,1,\dots,\left\lfloor C_{in}/2\right\rfloor-1\}$ (1)
where $F_{even},F_{odd}\in\mathbb{R}^{B\times(C_{in}/2)\times H\times W}$. Guided by the feature pyramid theory, convolutional kernels of differing scales ($\times3$ and $\times5$) are implemented through depth-wise separable convolutions on the decoupled subfeatures respectively, enhancing cross-scale feature completeness [23]. After multi-kernel processing, the decomposed features undergo:
@@
where $\alpha$ = 0.05 governs the negative signal retention ratio, enhancing robust-ness against adversarial samples. $\mathcal{W}_{\textsl{pw}}$ denotes the $\times1$ pointwise convolution.
@@ 2.2.2 Cross-Resolution Hierarchical Feature Fusion
[…]
The first layer of the encoder processes single-channel input $X_{a}\in\mathbb{R}^{1\times 256\times 256}$, preserving resolution while expanding the channel dimension to 3. Reflection padding(nn.ReflectionPad2d(3))extends it to $262\times 262$, mirroring boundaries to reduce artifacts.A $7\times 7$ convolution (a stride of 2) outputs a $3\times 256\times 256$ feature map, followed by BatchNorm and ReLU, yielding $f_{1}^{a}\in\mathbb{R}^{3\times 256\times 256}$. For the second layer, a $\times3$ convolution with stride=2 downsamples $f_{1}^{a}$ to $f_{2}^{a}\in\mathbb{R}^{144\times 128\times 128}$. After normalization and activation procedures, the RFFN integrates $f_{2}^{a}$ and $f_{1}^{b}$ via residual connections, generating the fused feature $f_{21}^{ab}\in\mathbb{R}^{144\times 128\times 128}$. The third layer follows this paradigm, further downsampling to $f_{3}^{ab}\in\mathbb{R}^{288\times 64\times 64}$ and performing secondary fusion with $f_{2}^{b}$, ultimately producing $f_{32}^{ab}\in\mathbb{R}^{288\times 64\times 64}$.
@@ 2.3 Bidirectional Mamba Residual Network
$\left\{\begin{aligned} \tilde{X}_{1}&=\text{Reshape}(X)\in\mathbb{R}^{B\times C\times L}\\ \tilde{X}_{1}&=S\cdot\hat{X}_{1}^{\top}\in\mathbb{R}^{B\times L\times C}\end{aligned}\right.L=H\times W$ (3)
@@
$\left\{\begin{aligned} Y_{1}&=\text{Mamba}_{1}(\text{LN}(\tilde{X}_{1}))\\ Y_{2}&=\text{Mamba}_{2}(\text{LN}(\tilde{X_{2}}))\end{aligned}\right.\quad\text{Mamba}(Z)=\sigma(W_{\text{conv}}*Z)\odot(W_{\text{proj}}\cdot\text{SSM}(Z))$ (4)
@@ 3.2 Results
[…]
@@
[…]